OpenStudy (anonymous):

Find the speed of a particle whose total energy is 50% greater than its rest energy.

OpenStudy (vincent-lyon.fr):

\(E^2=p^2c^2+m^2c^4\)

OpenStudy (anonymous):

\[E _{total}=\gamma mc^2\] \[E _{rest}=mc^2\] \[3/2E _{total}=E _{rest}\] \[3/2 \gamma mc^2 = mc^2\]\[3/2 \gamma = 1\] Solve for V

OpenStudy (vincent-lyon.fr):

Much better solution given by azolotor than me. Just line misprinted: \(E _{total}=3/2 \, E _{rest}\) \(\gamma = 3/2\)